# Correlated Proxies: A New Definition and Improved Mitigation for Reward Hacking

### v3

@@ A.1.1 Near-Optimal Base Policies
Suppose that $\tilde{R}$ is an $r$-correlated-proxy for the true reward function $R$, and let $\sigma_{\tilde{R}}$ and $\sigma_{R}$ be defined as in Definition 4.1. Furthermore, suppose that the base policy $\pi_{\text{base}}$} is near-optimal:
$J({\pi_{\text{base}}},R)\geq\max_{\pi^{*}}J(\pi^{*},R)-\epsilon\sigma_{R},$
@@
Then for any policy $\pi$ such that $\mu_{\pi}\ll\mu_{{\pi_{\text{base}}}}$, we have
$\frac{\max_{\pi^{*}}J(\pi^{*},R)-J(\pi,R)}{\sigma_{R}}\leq\epsilon-\frac{1}{r}\left(\frac{J(\pi,\tilde{R})-J({\pi_{\text{base}}},\tilde{R})}{\sigma_{\tilde{R}}}-\sqrt{(1-r^{2})\chi^{2}\left(\mu_{\pi}\|\mu_{\pi_{\text{base}}}\right)}\right).$ (10)
This result bounds the suboptimality gap—how close the learned policy is to optimal—in terms of the suboptimality gap of the base policy and the increase in the regularized proxy reward. The proof is straightforward and follows from Theorem 5.1.
@@
$L(\pi)=\frac{1}{r}\left(\frac{J(\pi,\tilde{R})-J({\pi_{\text{base}}},\tilde{R})}{\sigma_{\tilde{R}}}-\sqrt{(1-r^{2})\chi^{2}\left(\mu_{\pi}\|\mu_{\pi_{\text{base}}}\right)}\right)$ (11)
@@
$L(\pi)\leq\frac{1-\sqrt{1-r^{2}}}{r}\sqrt{\chi^{2}\left(\mu_{\pi}\|\mu_{\pi_{\text{base}}}\right)}.$ (12)
This shows that the lower bound can be at most a factor of $\frac{1-\sqrt{1-r^{2}}}{r}$ times the divergence between the learned and base policies’ occupancy measures. This factor is increasing in $r$ and asymptotes to $r/2$ as $r\to 0$.
@@
$L(\pi)=\frac{1}{r}\left(\mathbb{E}_{\pi}\left[\tilde{Z}(s,a)\right]-\sqrt{(1-r^{2})\chi^{2}\left(\mu_{\pi}\|\mu_{\pi_{\text{base}}}\right)}\right).$
@@
$\mathbb{E}_{\pi}\left[\tilde{Z}(s,a)\right]$ $=\mathbb{E}_{\pi}\left[\tilde{Z}(s,a)\right]-\mathbb{E}_{\pi_{\text{base}}}\left[\tilde{Z}(s,a)\right]$
$=\sum_{(s,a)\in\mathcal{S}\times\mathcal{A}}\left[\sqrt{\mu_{\pi_{\text{base}}}(s,a)}\tilde{Z}(s,a)\right]\left[\frac{\mu_{\pi}(s,a)-\mu_{\pi_{\text{base}}}(s,a)}{\sqrt{\mu_{\pi_{\text{base}}}(s,a)}}\right]$
$\leq\sqrt{\left(\sum_{(s,a)\in\mathcal{S}\times\mathcal{A}}\mu_{\pi_{\text{base}}}(s,a)\tilde{Z}(s,a)^{2}\right)\left(\sum_{(s,a)\in\mathcal{S}\times\mathcal{A}}\frac{(\mu_{\pi}(s,a)-\mu_{\pi_{\text{base}}}(s,a))^{2}}{\mu_{\pi_{\text{base}}}(s,a)}\right)}$
$=\sqrt{\mathbb{E}_{\mu_{\pi_{\text{base}}}}\left[\tilde{Z}(s,a)^{2}\right]\chi^{2}\left(\mu_{\pi}\|\mu_{\pi_{\text{base}}}\right)}$
$=\sqrt{\chi^{2}\left(\mu_{\pi}\|\mu_{\pi_{\text{base}}}\right)}.$
@@
$L(\pi)$ $\leq\frac{1}{r}\left(\sqrt{\chi^{2}\left(\mu_{\pi}\|\mu_{\pi_{\text{base}}}\right)}-\sqrt{(1-r^{2})\chi^{2}\left(\mu_{\pi}\|\mu_{\pi_{\text{base}}}\right)}\right)$
$=\frac{1-\sqrt{1-r^{2}}}{r}\sqrt{\chi^{2}\left(\mu_{\pi}\|\mu_{\pi_{\text{base}}}\right)},$
@@ A.1.3 Is the Lower Bound Optimizable?
While Theorem 5.1 shows that the increase in true reward over the base policy can be lower-bounded by the the proxy reward minus a regularization term, it is not clear if it is actually possible to increase the lower bound $L(\pi)$ as defined in (11). For example, if the base policy is already optimal with respect to the proxy reward, then clearly $L(\pi)\leq 0$. As another example, suppose it possible to improve $\pi_{\text{base}}$} with respect to both the true and proxy rewards, but only by visiting a state-action pair never visited by $\pi_{\text{base}}$}. In this case, it is also impossible to improve the lower bound in Theorem 5.1 while obeying the requirement that $\mu_{\pi}\ll\mu_{{\pi_{\text{base}}}}$.
@@
Fix any $r\in(0,1)$. Then there is an MDP with a true reward function $R$, a proxy reward $\tilde{R}$, and a base policy $\pi_{\text{base}}$} such that $\tilde{R}$ is an $r$-correlated proxy that can be improved upon in both true and proxy reward by a policy $\pi^{*}$:
$J(\pi^{*},R)$ $>J({\pi_{\text{base}}},R)$
@@
$\mu_{\pi^{*}}$ $\ll\mu_{\pi_{\text{base}}}.$
However, for any policy $\pi$ such that $\mu_{\pi}\ll\mu_{{\pi_{\text{base}}}}$,
$L(\pi)=\frac{1}{r}\left(\frac{J(\pi,\tilde{R})-J({\pi_{\text{base}}},\tilde{R})}{\sigma_{\tilde{R}}}-\sqrt{(1-r^{2})\chi^{2}\left(\mu_{\pi}\|\mu_{\pi_{\text{base}}}\right)}\right)\leq 0.$
That is, Lemma A.3 shows that there is an MDP where there exists a policy $\pi^{*}$ that improves on the base policy in both true reward and proxy reward, but it is still not possible to increase the lower bound $L(\pi)$ above zero. This suggests that it is difficult to specify general conditions under which $L(\pi)$ can exceed zero. Thus, we rely on the empirical evidence from our experiments to show that the lower bound is often optimizable.
@@
Furthermore, we define ${\pi_{\text{base}}}(a_{1}\mid s_{1})=1-r$ and ${\pi_{\text{base}}}(a_{2}\mid s_{1})=r$, and ${\pi_{\text{base}}}(a_{1}\mid s_{2})=1$. Based on this, simple algebra shows the following facts:
$\mu_{\pi_{\text{base}}}(s_{1},a_{1})$ $=\frac{1-r}{1+r}$ $\qquad\mu_{\pi_{\text{base}}}(s_{1},a_{2})$ $=\frac{r}{1+r}$
@@
$\mathbb{E}_{\mu_{{\pi_{\text{base}}}}}\left[R(s,a)\tilde{R}(s,a)\right]$ $=\frac{r}{1+r}.$
@@
$J(\pi^{*},R)$ $=\frac{1}{1+r}\sqrt{\frac{r}{1-r}}>0=J({\pi_{\text{base}}},R)$
@@
Now, consider any policy $\pi$ such that $\mu_{\pi}\ll\mu_{{\pi_{\text{base}}}}$. We can calculate the lower bound $L(\pi)$ (ignoring the prefactor of $\frac{1}{r}$) as
$J(\pi,\tilde{R})\sqrt{1+r}-\sqrt{(1-r^{2})\chi^{2}\left(\mu_{\pi}\|\mu_{\pi_{\text{base}}}\right)}.$ (13)
Since $\mu_{\pi}\ll\mu_{{\pi_{\text{base}}}}$, $\pi(a_{1}\mid s_{2})=1$, so it can only differ from $\pi_{\text{base}}$} in state $s_{1}$. Let $\delta=\pi(a_{1}\mid s_{1})-{\pi_{\text{base}}}(a_{1}\mid s_{1})$. Then, we can write the first term of (13) as
@@
Therefore, the $\chi^{2}$ divergence between $\mu_{\pi}$ and $\mu_{\pi_{\text{base}}}$ can be lower bounded as
$\chi^{2}\left(\mu_{\pi}\|\mu_{\pi_{\text{base}}}\right)\geq\frac{\left(\mu_{\pi}(s_{1},a_{1})-\mu_{\pi_{\text{base}}}(s_{1},a_{1})\right)^{2}}{\mu_{\pi_{\text{base}}}(s_{1},a_{1})}=\frac{\left(\frac{1-r+\delta}{1+r}-\frac{1-r}{1+r}\right)^{2}}{\frac{1-r}{1+r}}=\frac{\delta^{2}}{1-r^{2}}.$
@@
We define the base policy $\pi_{\text{base}}$} as follows:
${\pi_{\text{base}}}(a_{1}\mid s_{1})$ $=\frac{r^{2}}{2r^{2}-2r+1}\qquad$ ${\pi_{\text{base}}}(a_{2}\mid s_{1})$ $=\frac{(1-r)^{2}}{2r^{2}-2r+1}$
${\pi_{\text{base}}}(a_{1}\mid s_{2})$ $=1\qquad$ ${\pi_{\text{base}}}(a_{1}\mid s_{3})$ $=1.$
@@
$\mu_{\pi_{\text{base}}}(s_{1},a_{1})$ $=\frac{r^{2}}{r^{2}-r+1}$ $\qquad\mu_{\pi_{\text{base}}}(s_{1},a_{2})$ $=\frac{(1-r)^{2}}{r^{2}-r+1}$
@@
$\mathbb{E}_{\mu_{{\pi_{\text{base}}}}}\left[R(s,a)\tilde{R}(s,a)\right]$ $=\frac{r^{2}}{r^{2}-r+1}.$
@@
$J(\pi^{*},R)$ $=\frac{1-r}{r^{2}-r+1}\sqrt{\frac{1-r}{r}}>0=J({\pi_{\text{base}}},R)$
@@
Now, consider any policy $\pi$ such that $\mu_{\pi}\ll\mu_{{\pi_{\text{base}}}}$. We can calculate the lower bound $L(\pi)$ (again ignoring the prefactor of $\frac{1}{r}$) as
$J(\pi,\tilde{R})\sqrt{\frac{r^{2}-r+1}{r}}-\sqrt{(1-r^{2})\chi^{2}\left(\mu_{\pi}\|\mu_{\pi_{\text{base}}}\right)}.$ (14)
As in the previous case, let $\delta=\pi(a_{1}\mid s_{1})-{\pi_{\text{base}}}(a_{1}\mid s_{1})$. Then, we can write the first term of (14) as
@@
The $\chi^{2}$ divergence between $\mu_{\pi}$ and $\mu_{\pi_{\text{base}}}$ can be lower bounded as
$\chi^{2}\left(\mu_{\pi}\|\mu_{\pi_{\text{base}}}\right)\geq\frac{\left(\mu_{\pi}(s_{1},a_{1})-\mu_{\pi_{\text{base}}}(s_{1},a_{1})\right)^{2}}{\mu_{\pi_{\text{base}}}(s_{1},a_{1})}=\frac{\left(\frac{2r^{2}-2r+1}{r^{2}-r+1}\delta\right)^{2}}{\frac{r^{2}}{r^{2}-r+1}}=\delta^{2}\left(\frac{2r^{2}-2r+1}{r}\right)^{2}\frac{1}{r^{2}-r+1}.$
@@
Fix any $r\in(0,1)$. Then there is an MDP with a true reward function $R$, a proxy reward $\tilde{R}$, and a base policy $\pi_{\text{base}}$} such that $\tilde{R}$ is an $r$-correlated proxy and:
@@
We define the base policy $\pi_{\text{base}}$} as follows:
${\pi_{\text{base}}}(a_{1}\mid s_{1})={\pi_{\text{base}}}(a_{2}\mid s_{1})=1/2\qquad{\pi_{\text{base}}}(a_{1}\mid s_{2})=1\qquad{\pi_{\text{base}}}(a_{1}\mid s_{3})=1.$
@@
$\mathbb{E}_{\mu_{\pi_{\text{base}}}}\Big[R(s,a)\tilde{R}(s,a)\Big]$
@@
Since the rewards for both actions are identical at $s_{2}$ and $s_{3}$, a policy can only differ meaningfully from $\pi_{\text{base}}$} at $s_{1}$. Define
@@
$L(\pi_{\Delta})$ $=\frac{1}{r}\left(J(\pi_{\Delta},\tilde{R})-\sqrt{(1-r^{2})\chi^{2}\left(\mu_{\pi_{\Delta}}\|\mu_{\pi_{\text{base}}}\right)}\right)$

### v4

@@ A.1.1 Near-Optimal reference policies
Suppose that $\tilde{R}$ is an $r$-correlated-proxy for the true reward function $R$, and let $\sigma_{\tilde{R}}$ and $\sigma_{R}$ be defined as in Definition 4.1. Furthermore, suppose that the reference policy $\pi_{\text{ref}}$} is near-optimal:
$J({\pi_{\text{ref}}},R)\geq\max_{\pi^{*}}J(\pi^{*},R)-\epsilon\sigma_{R},$
@@
Then for any policy $\pi$ such that $\mu_{\pi}\ll\mu_{{\pi_{\text{ref}}}}$, we have
$\frac{\max_{\pi^{*}}J(\pi^{*},R)-J(\pi,R)}{\sigma_{R}}\leq\epsilon-\frac{1}{r}\left(\frac{J(\pi,\tilde{R})-J({\pi_{\text{ref}}},\tilde{R})}{\sigma_{\tilde{R}}}-\sqrt{(1-r^{2})\chi^{2}\left(\mu_{\pi}\|\mu_{\pi_{\text{ref}}}\right)}\right).$ (10)
This result bounds the suboptimality gap—how close the learned policy is to optimal—in terms of the suboptimality gap of the reference policy and the increase in the regularized proxy reward. The proof is straightforward and follows from Theorem 5.1.
@@
$L(\pi)=\frac{1}{r}\left(\frac{J(\pi,\tilde{R})-J({\pi_{\text{ref}}},\tilde{R})}{\sigma_{\tilde{R}}}-\sqrt{(1-r^{2})\chi^{2}\left(\mu_{\pi}\|\mu_{\pi_{\text{ref}}}\right)}\right)$ (11)
@@
$L(\pi)\leq\frac{1-\sqrt{1-r^{2}}}{r}\sqrt{\chi^{2}\left(\mu_{\pi}\|\mu_{\pi_{\text{ref}}}\right)}.$ (12)
This shows that the lower bound can be at most a factor of $\frac{1-\sqrt{1-r^{2}}}{r}$ times the divergence between the learned and reference policies’ occupancy measures. This factor is increasing in $r$ and asymptotes to $r/2$ as $r\to 0$.
@@
$L(\pi)=\frac{1}{r}\left(\mathbb{E}_{\pi}\left[\tilde{Z}(s,a)\right]-\sqrt{(1-r^{2})\chi^{2}\left(\mu_{\pi}\|\mu_{\pi_{\text{ref}}}\right)}\right).$
@@
$\mathbb{E}_{\pi}\left[\tilde{Z}(s,a)\right]$ $=\mathbb{E}_{\pi}\left[\tilde{Z}(s,a)\right]-\mathbb{E}_{\pi_{\text{ref}}}\left[\tilde{Z}(s,a)\right]$
$=\sum_{(s,a)\in\mathcal{S}\times\mathcal{A}}\left[\sqrt{\mu_{\pi_{\text{ref}}}(s,a)}\tilde{Z}(s,a)\right]\left[\frac{\mu_{\pi}(s,a)-\mu_{\pi_{\text{ref}}}(s,a)}{\sqrt{\mu_{\pi_{\text{ref}}}(s,a)}}\right]$
$\leq\sqrt{\left(\sum_{(s,a)\in\mathcal{S}\times\mathcal{A}}\mu_{\pi_{\text{ref}}}(s,a)\tilde{Z}(s,a)^{2}\right)\left(\sum_{(s,a)\in\mathcal{S}\times\mathcal{A}}\frac{(\mu_{\pi}(s,a)-\mu_{\pi_{\text{ref}}}(s,a))^{2}}{\mu_{\pi_{\text{ref}}}(s,a)}\right)}$
$=\sqrt{\mathbb{E}_{\mu_{\pi_{\text{ref}}}}\left[\tilde{Z}(s,a)^{2}\right]\chi^{2}\left(\mu_{\pi}\|\mu_{\pi_{\text{ref}}}\right)}$
$=\sqrt{\chi^{2}\left(\mu_{\pi}\|\mu_{\pi_{\text{ref}}}\right)}.$
@@
$L(\pi)$ $\leq\frac{1}{r}\left(\sqrt{\chi^{2}\left(\mu_{\pi}\|\mu_{\pi_{\text{ref}}}\right)}-\sqrt{(1-r^{2})\chi^{2}\left(\mu_{\pi}\|\mu_{\pi_{\text{ref}}}\right)}\right)$
$=\frac{1-\sqrt{1-r^{2}}}{r}\sqrt{\chi^{2}\left(\mu_{\pi}\|\mu_{\pi_{\text{ref}}}\right)},$
@@ A.1.3 Is the Lower Bound Optimizable?
While Theorem 5.1 shows that the increase in true reward over the reference policy can be lower-bounded by the the proxy reward minus a regularization term, it is not clear if it is actually possible to increase the lower bound $L(\pi)$ as defined in (11). For example, if the reference policy is already optimal with respect to the proxy reward, then clearly $L(\pi)\leq 0$. As another example, suppose it possible to improve $\pi_{\text{ref}}$} with respect to both the true and proxy rewards, but only by visiting a state-action pair never visited by $\pi_{\text{ref}}$}. In this case, it is also impossible to improve the lower bound in Theorem 5.1 while obeying the requirement that $\mu_{\pi}\ll\mu_{{\pi_{\text{ref}}}}$.
@@
Fix any $r\in(0,1)$. Then there is an MDP with a true reward function $R$, a proxy reward $\tilde{R}$, and a reference policy $\pi_{\text{ref}}$} such that $\tilde{R}$ is an $r$-correlated proxy that can be improved upon in both true and proxy reward by a policy $\pi^{*}$:
$J(\pi^{*},R)$ $>J({\pi_{\text{ref}}},R)$
@@
$\mu_{\pi^{*}}$ $\ll\mu_{\pi_{\text{ref}}}.$
However, for any policy $\pi$ such that $\mu_{\pi}\ll\mu_{{\pi_{\text{ref}}}}$,
$L(\pi)=\frac{1}{r}\left(\frac{J(\pi,\tilde{R})-J({\pi_{\text{ref}}},\tilde{R})}{\sigma_{\tilde{R}}}-\sqrt{(1-r^{2})\chi^{2}\left(\mu_{\pi}\|\mu_{\pi_{\text{ref}}}\right)}\right)\leq 0.$
That is, Lemma A.3 shows that there is an MDP where there exists a policy $\pi^{*}$ that improves on the reference policy in both true reward and proxy reward, but it is still not possible to increase the lower bound $L(\pi)$ above zero. This suggests that it is difficult to specify general conditions under which $L(\pi)$ can exceed zero. Thus, we rely on the empirical evidence from our experiments to show that the lower bound is often optimizable.
@@
Furthermore, we define ${\pi_{\text{ref}}}(a_{1}\mid s_{1})=1-r$ and ${\pi_{\text{ref}}}(a_{2}\mid s_{1})=r$, and ${\pi_{\text{ref}}}(a_{1}\mid s_{2})=1$. Based on this, simple algebra shows the following facts:
$\mu_{\pi_{\text{ref}}}(s_{1},a_{1})$ $=\frac{1-r}{1+r}$ $\qquad\mu_{\pi_{\text{ref}}}(s_{1},a_{2})$ $=\frac{r}{1+r}$
@@
$\mathbb{E}_{\mu_{{\pi_{\text{ref}}}}}\left[R(s,a)\tilde{R}(s,a)\right]$ $=\frac{r}{1+r}.$
@@
$J(\pi^{*},R)$ $=\frac{1}{1+r}\sqrt{\frac{r}{1-r}}>0=J({\pi_{\text{ref}}},R)$
@@
Now, consider any policy $\pi$ such that $\mu_{\pi}\ll\mu_{{\pi_{\text{ref}}}}$. We can calculate the lower bound $L(\pi)$ (ignoring the prefactor of $\frac{1}{r}$) as
$J(\pi,\tilde{R})\sqrt{1+r}-\sqrt{(1-r^{2})\chi^{2}\left(\mu_{\pi}\|\mu_{\pi_{\text{ref}}}\right)}.$ (13)
Since $\mu_{\pi}\ll\mu_{{\pi_{\text{ref}}}}$, $\pi(a_{1}\mid s_{2})=1$, so it can only differ from $\pi_{\text{ref}}$} in state $s_{1}$. Let $\delta=\pi(a_{1}\mid s_{1})-{\pi_{\text{ref}}}(a_{1}\mid s_{1})$. Then, we can write the first term of (13) as
@@
Therefore, the $\chi^{2}$ divergence between $\mu_{\pi}$ and $\mu_{\pi_{\text{ref}}}$ can be lower bounded as
$\chi^{2}\left(\mu_{\pi}\|\mu_{\pi_{\text{ref}}}\right)\geq\frac{\left(\mu_{\pi}(s_{1},a_{1})-\mu_{\pi_{\text{ref}}}(s_{1},a_{1})\right)^{2}}{\mu_{\pi_{\text{ref}}}(s_{1},a_{1})}=\frac{\left(\frac{1-r+\delta}{1+r}-\frac{1-r}{1+r}\right)^{2}}{\frac{1-r}{1+r}}=\frac{\delta^{2}}{1-r^{2}}.$
@@
We define the reference policy $\pi_{\text{ref}}$} as follows:
${\pi_{\text{ref}}}(a_{1}\mid s_{1})$ $=\frac{r^{2}}{2r^{2}-2r+1}\qquad$ ${\pi_{\text{ref}}}(a_{2}\mid s_{1})$ $=\frac{(1-r)^{2}}{2r^{2}-2r+1}$
${\pi_{\text{ref}}}(a_{1}\mid s_{2})$ $=1\qquad$ ${\pi_{\text{ref}}}(a_{1}\mid s_{3})$ $=1.$
@@
$\mu_{\pi_{\text{ref}}}(s_{1},a_{1})$ $=\frac{r^{2}}{r^{2}-r+1}$ $\qquad\mu_{\pi_{\text{ref}}}(s_{1},a_{2})$ $=\frac{(1-r)^{2}}{r^{2}-r+1}$
@@
$\mathbb{E}_{\mu_{{\pi_{\text{ref}}}}}\left[R(s,a)\tilde{R}(s,a)\right]$ $=\frac{r^{2}}{r^{2}-r+1}.$
@@
$J(\pi^{*},R)$ $=\frac{1-r}{r^{2}-r+1}\sqrt{\frac{1-r}{r}}>0=J({\pi_{\text{ref}}},R)$
@@
Now, consider any policy $\pi$ such that $\mu_{\pi}\ll\mu_{{\pi_{\text{ref}}}}$. We can calculate the lower bound $L(\pi)$ (again ignoring the prefactor of $\frac{1}{r}$) as
$J(\pi,\tilde{R})\sqrt{\frac{r^{2}-r+1}{r}}-\sqrt{(1-r^{2})\chi^{2}\left(\mu_{\pi}\|\mu_{\pi_{\text{ref}}}\right)}.$ (14)
As in the previous case, let $\delta=\pi(a_{1}\mid s_{1})-{\pi_{\text{ref}}}(a_{1}\mid s_{1})$. Then, we can write the first term of (14) as
@@
The $\chi^{2}$ divergence between $\mu_{\pi}$ and $\mu_{\pi_{\text{ref}}}$ can be lower bounded as
$\chi^{2}\left(\mu_{\pi}\|\mu_{\pi_{\text{ref}}}\right)\geq\frac{\left(\mu_{\pi}(s_{1},a_{1})-\mu_{\pi_{\text{ref}}}(s_{1},a_{1})\right)^{2}}{\mu_{\pi_{\text{ref}}}(s_{1},a_{1})}=\frac{\left(\frac{2r^{2}-2r+1}{r^{2}-r+1}\delta\right)^{2}}{\frac{r^{2}}{r^{2}-r+1}}=\delta^{2}\left(\frac{2r^{2}-2r+1}{r}\right)^{2}\frac{1}{r^{2}-r+1}.$
@@
Fix any $r\in(0,1)$. Then there is an MDP with a true reward function $R$, a proxy reward $\tilde{R}$, and a reference policy $\pi_{\text{ref}}$} such that $\tilde{R}$ is an $r$-correlated proxy and:
@@
We define the reference policy $\pi_{\text{ref}}$} as follows:
${\pi_{\text{ref}}}(a_{1}\mid s_{1})={\pi_{\text{ref}}}(a_{2}\mid s_{1})=1/2\qquad{\pi_{\text{ref}}}(a_{1}\mid s_{2})=1\qquad{\pi_{\text{ref}}}(a_{1}\mid s_{3})=1.$
@@
$\mathbb{E}_{\mu_{\pi_{\text{ref}}}}\Big[R(s,a)\tilde{R}(s,a)\Big]$
@@
Since the rewards for both actions are identical at $s_{2}$ and $s_{3}$, a policy can only differ meaningfully from $\pi_{\text{ref}}$} at $s_{1}$. Define
@@
$L(\pi_{\Delta})$ $=\frac{1}{r}\left(J(\pi_{\Delta},\tilde{R})-\sqrt{(1-r^{2})\chi^{2}\left(\mu_{\pi_{\Delta}}\|\mu_{\pi_{\text{ref}}}\right)}\right)$